# DEER: A Comprehensive and Reliable Benchmark for Deep-Research Expert Reports

### v1

@@ Abstract
As large language models (LLMs) advance, deep research systems can generate expert-level reports via multi-step reasoning and evidence-based synthesis, but evaluating such reports remains challenging. Existing benchmarks often lack systematic criteria for expert reporting, evaluations that rely heavily on LLM judges can fail to capture issues that require expert judgment, and source verification typically covers only a limited subset of explicitly cited statements rather than report-wide factual reliability. We introduce DEER, a benchmark for evaluating expert-level deep research reports. DEER comprises 50 report-writing tasks spanning 13 domains and an expert-grounded evaluation taxonomy (7 dimensions, 25 sub-dimension) operationalized into 130 fine-grained rubric items. DEER further provides task-specific expert guidance to help LLM judges assess expert-level report quality more consistently. Complementing rubric-based assessment, we propose a document-level fact-checking architecture that extracts and verifies all claims across the entire report, including both cited and uncited ones, and quantifies external-evidence quality. DEER correlates closely with human expert judgments and yields interpretable diagnostics of system strengths and weaknesses.
DEER: A Comprehensive and Reliable Benchmark
for Deep-Research Expert Reports
@@ 1 Introduction
Driven by the rapid advancement of large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry openai2025deepresearch; google2025gemini; anthropic2025claude; qwen3technicalreport; webthinker; manusearch; webweaver. Unlike simple information retrieval, deep research systems tackle complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results. This enables them to integrate information from diverse sources and analyze different perspectives in a balanced manner, producing reliable, evidence-based research reports deepresearch_survey; deepresearch_survey2; characterizing. These advanced capabilities enable deep research systems to achieve strong performance even on challenging benchmark tasks.
@@
Figure 1: Deep Research System Performance Comparison. Shows the performance by type for 5 major models in the proposed benchmark.
@@
Early evaluation of deep research systems relied on complex reasoning benchmarks gpqa; hle, which indirectly assessed information-gathering, hypothesis testing, and multi-step reasoning through performance on these tasks. Subsequently, deep web search QA benchmarks were developed to directly assess systems’ web browsing and information retrieval abilities—core capabilities of deep research—by measuring multi-step search, information integration, and answer derivation browsecomp; frames; gaia; browsecomp_plus; mind2web2. More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, going beyond simple short-answer evaluation webthinker; deepbench; deepconsult; deepgym.
@@
Despite these advancements, current methodologies for evaluating deep research systems still face key limitations in assessing expert reports. First, while prior work has proposed high-level dimensions for report evaluation, these frameworks have yet to comprehensively specify the diverse elements required for expert-level reports. In particular, many frameworks deepbench; DeepScholar-Bench; LiveResearchBench; DeepResearchArena rely on LLMs—rather than human experts—to instantiate detailed sub-dimension and concrete evaluation items, which can lead to inconsistencies and reduced reliability. Second, even with a fixed rubric, evaluations that rely on LLM judges can fail to capture issues that require expert judgment. Third, current approaches to source verification are typically limited to sections with explicit citation markers, leaving report-wide claim reliability insufficiently examined. Together, these limitations hinder comprehensive and reliable evaluation of deep research systems.
@@
To address these limitations, we propose the Deep rEsearch Expert Report benchmark (DEER), consisting of 50 report-generation tasks spanning 13 domains, for comprehensive and reliable evaluation. DEER analyzes a wide range of expert report assessment criteria and consolidates them through expert consensus into a Deep Research Report Evaluation Taxonomy with 7 dimensions and 25 sub-dimension. Based on this taxonomy, DEER evaluates each report along two complementary components: (i) report-quality dimensions that require holistic document-level judgment and (ii) external-information dimensions that can be assessed at the claim level via fact verification. For report quality, we operationalize the taxonomy into a fixed set of 130 fine-grained rubric items and supplement it with report-specific Expert Evaluation Guidance authored by domain experts to improve the consistency and validity of LLM-judge-based scoring. For external information, we introduce a fact-verification module that checks all cited and uncited claims across the report and produces quantitative measures of external-evidence quality. Finally, DEER combines these metric-based signals with rubric-based LLM evaluations to yield an integrated, multi-dimensional assessment of each expert report. Figure 1 presents an overview of dimension-level performance for major deep research systems evaluated on DEER.
The main contributions of this study are as follows:
@@
  • We present DEER, a systematic and interpretable benchmark for deep research, built on a hierarchical evaluation taxonomy.
@@
  • We instantiate this taxonomy into a fixed set of 130 fine-grained rubric items and supplement it with report-specific Expert Evaluation Guidance, improving the consistency and reliability of LLM-judge-based evaluation.
@@
  • We propose a fact verification architecture that backtracks inter-claim dependency relations to retrieve and verify relevant citations, enabling more complete report-wide evaluation of all claims.
@@ 2 Related Works
2.1 LLM-based Evaluation Automation
With the proliferation of Large Language Models, various approaches have been proposed to automate human-centric evaluation tasks using LLMs. Most notably, the LLM-as-a-Judge method, which directly uses LLMs as evaluators, has been widely researched. JudgeLM zhu2023judgelm and G-Eval liu2023g demonstrate that models can score generated responses based on multiple criteria such as accuracy, consistency, and logic. In addition, the scope of automated LLM evaluation is expanding to pairwise comparison-based evaluations such as Chatbot Arena chiang2024chatbot, CheckList-style robustness and bias checks checkeval, and various automated scoring frameworks. However, these approaches often struggle to control bias or hallucinations in the LLM’s own judgments and do not adequately address evaluations that require complex evidence structures, such as long-form reports.
2.2 Deep Research and High-Difficulty QA Evaluation
Recently, Deep Research benchmarks have emerged that go beyond simple knowledge recall to evaluate complex reasoning, web browsing, and information integration capabilities. High-difficulty QA benchmarks such as HLE hle, GPQA gpqa, and FRAMES frames reveal the limitations of existing models by requiring long-context reasoning, expert-level thinking, and information synthesis skills. Furthermore, ManuSearch manusearch and Mind2Web-2 mind2web2 employ multi-agent web browsing and agent-as-a-judge architectures to evaluate models’ ability to replicate real-world research processes. Additionally, DeepResearchBench deepbench and WebThinker Eval webthinker present Deep Research report-centric evaluations that assess literature-based report generation and multi-document reasoning, classifying them as measuring "research capability" beyond single-answer QA. However, while these benchmarks evaluate correctness, they cannot precisely verify the validity of evidence, citations, and arguments within the report, focusing more on whether the correct answer is reached than on factuality.
2.3 Claim Verification
In automated claim verification, various approaches have been proposed to mitigate LLM hallucinations. TrustGPT huang2023trustgpt attempted to improve factuality based on RAG, and InFact rothermel-etal-2024-infact established a structural verification system by decomposing claim verification into a 6-step pipeline. Additionally, MAPLE zeng2024maple, FactDetect jafari2024robustclaimverificationfact, and ClaimCheck putta-etal-2025-claimcheck implemented sophisticated claim-level fact-checking by combining claim type classification, evidence extraction, and verification automation. Furthermore, research analyzing citation quality in detail, such as ALiiCE xu2024aliiceevaluatingpositionalfinegrained and CiteEval xu2025citeevalprincipledrivencitationevaluation, which evaluate citation accuracy and evidence alignment in academic texts, is increasing. However, most approaches focus on verifying explicit citations or single claims and fail to address issues such as implicit evidence, contextual citation inheritance, and Fair Use that appear in long-form reports.
@@ 3 Data Construction
To construct an evaluation dataset that reflects real-world usage, we analyzed 5,045 user queries collected from a deep research system and estimated the resulting topic distribution. For topic classification, we used the taxonomy of weborganizer, following prior work deepbench.
We use Humanity’s Last Exam (HLE) hle as a source of seed questions because its expert-written, high-difficulty, multi-disciplinary questions align with the expert-level topics addressed in deep research reports. To match our target topic distribution, we mapped our topic domains to the 13 HLE subject domains and sampled 50 seed questions accordingly.
Because HLE items are posed in a QA format, they require reformulation before they can serve as deep research report-generation tasks. We therefore asked domain experts (with at least a master’s degree or equivalent expertise in the relevant field) to review each item’s question, answer, and rationale to identify the underlying concepts, theories, and phenomena, and to reformulate the items as report- or paper-style topics for report generation. Each reformulated prompt was drafted by one domain expert and cross-reviewed by another expert from the same field, with iterative revisions as needed. During reformulation, we removed answer-revealing elements (e.g., factual conclusions, proofs, or specific answers) and instead provided only high-level writing directions, such as the intended scope of analysis and comparative perspectives, so that the model must develop the reasoning and narrative on its own. This reformulation shifts the task from reproducing a short answer to generating a report that requires expert-level analysis, reasoning, and interpretation. Detailed content is presented in Appendix B.
@@ 4.1 Overview
Figure 2: Overview of the DEER evaluation framework. (a) Research question and expert guidance generation from real-world deep research queries. (b) Construction of the Deep Research Evaluation Taxonomy consisting of 7 dimensions, 26 criteria, and 130 granular rubrics. (c) The DEER evaluation pipeline, integrating expert-guided LLM-as-a-judge scoring with claim extraction and information verification to assess deep research reports.
To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis characterizing; deepresearch_survey; deepresearch_survey2. Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy with 7 major dimensions (§ 4.2), which we group into 5 report-quality dimensions and 2 external-information dimensions. Report-quality dimensions primarily assess synthesis and presentation via holistic, document-level judgment, whereas external-information dimensions assess external evidence acquisition and use via claim-level verification.
Based on these differences in evaluation characteristics, we propose a hybrid architecture that integrates methodologies tailored to each component, as illustrated in Figure 2. Expert-Guided Report Quality Assessment (§ 4.3) uses an LLM-as-a-judge approach, combining a fixed set of granular evaluation items with report-specific Expert Evaluation Guidance authored by domain experts, to assess report quality against expert-designed criteria. Information Verification (§ 4.4) performs metric-based evaluation by automatically extracting claims and citations and verifying claims against external sources, producing quantitative metrics of external evidence sufficiency and integrity. Scores from the rubric items and verification metrics are aggregated into 7 dimension scores, which are then combined into an overall report score.
Major Dimensions Sub-criteria
Request Completeness Request Completeness, Scope Boundary, Value Validity
Evidence Validity Numeric Accuracy, Logical Support
Structure Consistency Structure Intro, Structure Body, Structure Conclusion, Structure Section
Narration Style Report Form, Writing Quality, Paragraph, Reader Friendliness
Ethics Compliance Sensitive Issues, Safety Impact, Perspective Balance
Information Sufficiency Source Support, Information, Citations, References
Information Integrity Claim Factuality, Citation Validity, Reference Accuracy, Reference Quality, Reference Diversity
Table 1: Deep Research Report Evaluation Taxonomy: 7 Major Dimensions and 25 Sub-criteria
@@ 4.2 Deep Research Report Evaluation Taxonomy
In the absence of systematic evaluation criteria for deep research reports, we construct an evaluation taxonomy by synthesizing expert report assessment standards from multiple fields. We draw on internationally established standards spanning 20 domains of expertise, including systematic research reporting guidelines, technical and professional report writing and evaluation guidelines, and academic publishing norms, for a total of 80 standards. A panel of experts with experience in deep research system development and academic reviewing iteratively analyzes and consolidates these standards to derive report-quality dimensions. We then introduce complementary external-information dimensions that reflect deep research–specific characteristics of external evidence use and verification, resulting in 7 major dimensions and 25 detailed criteria in total (Table 1). We validate the taxonomy through independent review by 10 experts from diverse domains. This hierarchical taxonomy organizes diverse quality elements into structured evaluation dimensions and sub-dimension, enabling systematic, multi-dimensional evaluation. It facilitates diagnosis of report strengths and weaknesses and guides systematic improvement. Full descriptions, validation procedures, and the mapping from source standards to our criteria are provided in Appendix C.
@@ 4.3 Expert-Guided Report Quality Assessment
Existing evaluation methods have two limitations. First, existing methods apply broad, high-level criteria to long-form reports, giving LLM judges substantial discretion over what to focus on. As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) even focus on non-critical or superficial sub-aspects, increasing variance across judges and reducing evaluation consistency webthinker; deepconsult; deepgym. Second, even with a clear evaluation focus, assessing correctness and completeness often requires domain knowledge, which LLM judges may lack. As a result, they may miss subtle but important issues such as non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies deepbench. This study addresses these issues by (1) making the evaluation well-specified through a fixed set of fine-grained rubric items and (2) providing task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, helping LLM judges identify hard-to-detect errors and omissions.
@@
Granular Rubric Design. Broad evaluation criteria can lead to inconsistent LLM-judge-based scoring. Recent work addresses this by decomposing broad criteria into granular rubrics checkeval; rocketeval; expertlongbench. For deep research report evaluation, deepbench follows a similar direction by using an LLM to generate task-specific evaluation criteria. While this narrows the evaluation focus, it raises two concerns: whether the generated criteria are sufficiently complete and capture what matters most in expert reports, and whether dynamically changing criteria enable interpretable, comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
@@
This study uses a fixed, expert-designed rubric to ensure reliable and interpretable evaluation. Building on the taxonomy in Table 1, a panel of experts operationalizes the 25 sub-dimensions into 40 evaluation criteria and translates each criterion into concrete, checkable rubric items. Rubric items are organized into two aspects: coverage (whether the required components are present at all relevant points in the report and addressed in full whenever they occur) and quality (the degree to which the targeted components are executed to a high standard in the report). This process yields 130 scorable rubric items in total. [^1] The 40 criteria were validated through independent review by 10 experts following the procedure in § 4.2. The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix E for the rubric structure and examples).
Expert Evaluation Guidance. Even with a well-specified, granular rubric, scoring individual items can still require judgment informed by deep domain knowledge, especially for specialized and complex tasks—for example, to assess whether the report’s reasoning is sound, its calculations are correct, and its coverage and depth are adequate for the topic. In such cases, evaluators may miss subtle domain-specific errors or omissions that human experts would catch. To mitigate this, we provide task-specific expert evaluation guidance that enumerates the required content elements and expert expectations for each task, presented as concrete, checkable statements that evaluators can verify.
For each task, an expert with at least a master’s degree in the relevant field was assigned to detail the key concepts, theories, data, and analytical perspectives that the report must cover from an expert perspective. The expert first modified the original query to naturally include the essential content of the topic (Section 3), and then wrote the Expert Evaluation Guidance based on the modified query, specifying the content elements that must be included. This approach is based on the logic that if a report responds appropriately to the modified query, it should naturally contain the key content of the topic, so the omission of key content can be clearly judged as a quality issue. Each written Guidance was verified through cross-review by another expert in the same field. See Appendix B for detailed methodology.

### v2

# DEER: A Benchmark for Evaluating Deep Research Agents on Expert Report Generation

@@ Abstract
As large language models advance, deep research systems capable of generating expert-level reports through multi-step reasoning and evidence-based synthesis are emerging. However, evaluating such reports remains challenging. Existing benchmarks often lack systematic evaluation criteria, rely heavily on LLM-based judges that may miss issues requiring expert judgment, and verify only a limited subset of explicitly cited statements rather than report-wide factual reliability. To address these limitations, we introduce DEER, a benchmark for evaluating expert-level deep research reports. DEER comprises 50 report-writing tasks spanning 13 domains, along with an expert-grounded evaluation taxonomy with seven dimensions and 25 subdimensions, operationalized into 101 fine-grained rubric items. To improve evaluation consistency, DEER provides task-specific Expert Evaluation Guidance to support LLM-based judging. Complementing rubric-based assessment, we propose a document-level fact-checking architecture that verifies both cited and uncited claims and quantifies the quality and reliability of the supporting evidence. Experimental results show that DEER exhibits strong correlation with human expert judgments and yields interpretable diagnostics of system strengths and weaknesses.
DEER: A Benchmark for Evaluating Deep Research Agents
on Expert Report Generation
@@ 1 Introduction
Figure 1: Deep Research System Performance Comparison. Results for five systems on the proposed benchmark.
Driven by rapid advances in large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry OpenAI (2025); Google (2025); Anthropic (2025); Yang et al. (2025); Li et al. (2025b); Huang et al. (2025); Li et al. (2025c). Unlike traditional information retrieval systems, these systems address complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results. Through this process, they integrate information from diverse sources and synthesize multiple perspectives to produce reliable, evidence-based research reports Xu and Peng (2025); Zhang et al. (2025); Java et al. (2025). As a result, deep research systems demonstrate strong performance even on challenging benchmark tasks.
@@
Early evaluations of deep research systems relied primarily on complex reasoning benchmarks Rein et al. (2023); Phan et al. (2025), which indirectly assessed information-gathering, hypothesis testing, and multi-step reasoning through task performance. Subsequently, deep web search QA benchmarks were introduced to more directly measure systems’ web browsing and information retrieval abilities—core capabilities of deep research—by evaluating multi-step search, information integration, and answer derivation Wei et al. (2025a); Krishna et al. (2025); Mialon et al. (2023); Chen et al. (2025); Gou et al. (2025). More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, moving beyond simple short-answer-based evaluation Li et al. (2025b); Du et al. (2025); Consult (2025); Coelho et al. (2025).
@@
Despite these advancements, existing methodologies for evaluating deep research systems continue to face important limitations when applied to expert-level reports. First, although prior work has proposed high-level dimensions for report evaluation, these frameworks do not comprehensively specify the diverse elements required for expert reporting. In particular, many approaches Du et al. (2025); Patel et al. (2025); Wang et al. (2025); Wan et al. (2025) rely on LLMs rather than human experts to define detailed subdimensions and concrete evaluation items, which can introduce inconsistencies and reduce evaluation reliability. Second, even when a fixed rubric is provided, evaluations that rely on LLM judges may fail to identify issues that require expert judgment. Third, current approaches to source verification are typically restricted to sections with explicit citation markers, leaving the factual reliability of claims across the full report insufficiently examined. Together, these limitations hinder comprehensive and reliable evaluation of deep research systems.
@@
To address these limitations, we propose DEER (the DEep research Expert Report benchmark), which provides comprehensive and reliable evaluation of deep research reports through 50 report-generation tasks spanning 13 domains. DEER analyzes a broad range of expert report assessment criteria and consolidates them via expert consensus into a Deep Research Report Evaluation Taxonomy comprising 7 dimensions and 25 subdimensions. Based on this taxonomy, DEER evaluates each report across two complementary components: (i) report-quality dimensions that require holistic, document-level judgment, and (ii) external-information dimensions that can be assessed at the claim level by fact verification. For report quality, we operationalize the taxonomy into a fixed set of 101 fine-grained rubric items and supplement these with report-specific Expert Evaluation Guidance authored by domain experts to improve the consistency and validity of LLM-based scoring. For external information, we introduce a fact-verification module that examines both cited and uncited claims across the report and produces quantitative measures of evidence quality. Finally, DEER integrates rubric-based LLM evaluations with the resulting metric-based signals to yield a unified, multi-dimensional assessment of each expert report. Figure 1 presents an overview of dimension-level performance for major deep research systems evaluated using DEER.
The main contributions of this work are as follows:
@@
  • We present DEER, a systematic and interpretable benchmark for evaluating deep research reports, grounded in a hierarchical evaluation taxonomy.
@@
  • We operationalize this taxonomy into a fixed set of 101 fine-grained rubric items and provide report-specific Expert Evaluation Guidance, improving the consistency and reliability of LLM-based evaluation.
@@
  • We propose a report-level fact verification architecture that backtracks inter-claim dependencies to retrieve and verify relevant citations, enabling more complete evaluation of factual reliability across full reports.
Figure 2: Overview of the DEER evaluation framework. (a) Research question and expert guidance generation from real-world deep research queries. (b) Construction of the Deep Research Evaluation Taxonomy consisting of 7 dimensions, 25 sub-dimensions, and 101 granular rubrics. (c) The DEER evaluation pipeline, integrating expert-guided LLM-as-a-judge scoring with claim extraction and information verification to assess deep research reports.
@@ 2 Related Works
With the proliferation of high-performing LLMs, LLM-as-a-Judge approaches—using LLMs as evaluators—have been widely proposed and studied Zheng et al. (2023); Liu et al. (2023b); Chiang et al. (2024); Kim et al. (2024a, b). In this line of work, early evaluations for deep research primarily focused on how well models solve expert-level, high-difficulty questions Phan et al. (2025); Rein et al. (2023). Subsequently, to directly assess a core property of deep research—accessing and leveraging external information from the open web—benchmarks were proposed to measure models’ ability to browse/search the web, synthesize the required information, and construct answers grounded in that evidence Gou et al. (2025); Wei et al. (2025a); Huang et al. (2025). More recently, beyond simple short-answer evaluation, studies have sought to evaluate the ability to generate long-form reports that require expert-level analysis, reasoning, and interpretation by integrating information from multiple sources Consult (2025); Coelho et al. (2025); Du et al. (2025).
Despite this progress, existing report-evaluation methods are not based on expert-defined, fine-grained criteria. Some approaches rely on only a few coarse axes, leaving judgments largely to the evaluator LLM’s implicit standards Consult (2025); Coelho et al. (2025). Even when rubrics are subdivided, their specification and application can still depend substantially on the evaluator LLM Du et al. (2025), making it unclear whether the resulting scores correctly reflect report quality or align with expert assessment. In addition, source verification is often omitted Consult (2025); Coelho et al. (2025) or limited to a subset of cited claims Du et al. (2025), which may be insufficient to assess report-level factuality. To address these limitations, we propose an evaluation framework that (i) scores long-form research reports along multiple dimensions using expert-systematized, fine-grained criteria, and (ii) performs systematic source verification for claims throughout the report. A detailed comparison with prior work is provided in Appendix A.
@@ 3 Data Construction
To construct an evaluation dataset that reflects real-world usage, we analyze 5,842 in-house user queries collected from an internal deep research system and derive a topic distribution. For topic classification, we adopt the taxonomy of Wettig et al. (2025), consistent with prior work Du et al. (2025).
We use Humanity’s Last Exam (HLE) Phan et al. (2025) as the source of seed questions because it provides expert-written, high-difficulty, multidisciplinary questions that align well with the expert-level topics addressed in deep research reports. To match our target topic distribution, we map our topics to the 13 HLE subject domains and sample 50 seed questions accordingly.
Because HLE items are posed in a QA format, they need to be reformulated before they can be used as deep research report-generation queries. We therefore ask domain experts (each holding at least a master’s degree or possessing equivalent expertise in the relevant field) to review the original questions, answers, and rationales to identify the underlying concepts, theories, and phenomena, and to reformulate each item as a report- or paper-style prompt. Each reformulated prompt is drafted by one domain expert and cross-reviewed by another expert from the same field, with iterative revisions conducted as needed. During this process, we remove answer-revealing elements (e.g., factual conclusions, proofs, or specific answers) and retain only high-level writing directions, such as the intended scope of analysis and comparative perspectives. As a result, models are required to develop the reasoning and narrative independently. This reformulation shifts the task from producing a short answer to generating a report that requires expert-level analysis, reasoning, and interpretation. Detailed examples and procedures are provided in Appendix B.
@@ 4.1 Overview
To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis Java et al. (2025); Xu and Peng (2025); Zhang et al. (2025). Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy comprising seven major dimensions (§ 4.2), grouped into five report-quality dimensions and two external-information dimensions. The report-quality dimensions focus on synthesis and presentation through holistic, document-level judgment, whereas the external-information dimensions assess how external evidence is acquired and used via claim-level verification.
Given these differences in evaluation characteristics, we propose a hybrid evaluation architecture that integrates methodologies tailored to each component, as illustrated in Figure 2. Expert-Guided Report Quality Assessment (§ 4.3) adopts an LLM-as-a-judge approach, combining a fixed set of granular rubric items with report-specific Expert Evaluation Guidance authored by domain experts to assess report quality against expert-designed criteria. Information Verification (§ 4.4) performs metric-based evaluation by automatically extracting claims and citations, then verifying claims against external sources, yielding quantitative measures of the sufficiency and integrity of external evidence. Scores from the rubric-based assessment and information-verification metrics are aggregated into seven dimension-level scores, which are then combined into an overall report score.
@@ 4.2 Deep Research Report Evaluation Taxonomy
In the absence of systematic evaluation criteria for deep research reports, we construct an evaluation taxonomy by synthesizing expert report assessment standards from multiple fields. We draw on 80 established standards across 20 domains of expertise, including systematic research reporting guidelines, technical and professional report-writing and evaluation guidelines, and academic publishing norms. A panel of experts with experience in deep research system development and academic reviewing iteratively analyzes and consolidates these standards to derive report-quality dimensions. We then introduce complementary external-information dimensions that reflect deep research-specific characteristics of external evidence use and verification, resulting in a taxonomy comprising seven major dimensions and 25 detailed sub-dimensions in total (Table 9). We validate the taxonomy through independent review by 10 experts from diverse domains. This hierarchical taxonomy organizes diverse quality elements into structured evaluation dimensions and subdimensions, enabling systematic, multi-dimensional assessment of deep research reports. It supports diagnosis of report strengths and weaknesses and provides a foundation for targeted improvement. Full descriptions, validation procedures, and the mapping from source standards to our criteria are provided in Appendix C.
@@ 4.3 Expert-Guided Report Quality Assessment
Existing evaluation methods exhibit two key limitations when applied to long-form expert reports. First, they typically apply broad, high-level criteria, granting LLM judges substantial discretion over what aspects of a report to attend to. As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) focus on non-critical or superficial elements, leading to increased variance across judges and reduced evaluation consistency Li et al. (2025b); Consult (2025); Coelho et al. (2025). Second, even when the evaluation focus is clearly defined, assessing correctness and completeness often requires domain expertise that LLM judges may lack. Consequently, they may miss subtle but important issues, including non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies Du et al. (2025). To address these limitations, this study (1) makes evaluation criteria explicit through a fixed set of fine-grained rubric items and (2) provides task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, enabling LLM judges to identify hard-to-detect errors and omissions.
@@
Granular Rubric Design. Broad evaluation criteria can lead to inconsistent LLM-based scoring. Recent work addresses this issue by decomposing high-level criteria into more granular rubrics Lee et al. (2025); Wei et al. (2025b); Ruan et al. (2025). For deep research report evaluation, Du et al. (2025) follows a related approach by using an LLM to generate task-specific evaluation criteria. While this strategy narrows the evaluation focus, it raises two important concerns: whether the generated criteria are sufficiently comprehensive and capture what matters most in expert reports, and whether dynamically varying criteria enable interpretable and comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
@@
This study uses a fixed, expert-designed rubric to ensure reliable and interpretable evaluation. Building on the taxonomy in Table 9, a panel of experts operationalizes the 25 sub-dimensions into 46 evaluation criteria and translates each criterion into concrete, checkable rubric items. The items are organized into two aspects: coverage, whether required components are present and fully addressed wherever they occur in the report; and quality, the degree to which the targeted components are executed to a high standard. This process yields 101 scorable rubric items in total. [^1] The 46 criteria were validated through independent review by 10 experts following the procedure in § 4.2. The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix E for the rubric structure and examples).
Expert Evaluation Guidance. Even with a well-specified, fine-grained rubric, evaluating expert-level reports often requires substantial domain knowledge. In such settings, non-expert evaluators—including LLM judges—may fail to detect subtle domain-specific errors or omissions that subject-matter experts would identify. To mitigate this risk, we introduce task-specific Expert Evaluation Guidance, which provides domain-grounded context and reference points to support consistent and informed evaluation.
Expert Evaluation Guidance enumerates the required content elements and expert expectations for each task as concrete, verifiable statements that can be checked directly against the report. The guidance for each task is produced using the same expert drafting and cross-review procedure as the query reformulation process described in Section 3. Specifically, a domain expert reformulates each HLE item into a report-style prompt with explicit requirements (Section 3). The expert then derives the guidance by identifying the substantive content an adequate expert report should cover under the task prompt. Each required element is expressed as a concrete, verifiable statement that can be checked against the report. The resulting guidance is subsequently cross-reviewed by another expert from the same field to identify omissions, redundancies, or ambiguities, and revised as needed. Through this process, Expert Evaluation Guidance complements the fixed rubric by anchoring evaluation in task-specific expert knowledge while maintaining consistency and interpretability across tasks. Appendix B provides detailed guidance construction procedures and illustrative examples.
Model Req. Ful. Analyt. Sound. Struct. Coh. Format & Style Info. Int. Info. Suff. Ethics Mean
General LLMs
Qwen3-235B (fast) 4.51 5.02 6.09 7.49 1.24 4.20 7.19 5.11
Gemini 2.5 Flash (fast) 4.64 5.33 6.55 7.85 1.30 3.99 7.52 5.31
Claude Opus 4.5 (fast) 4.94 5.48 6.54 7.99 2.29 4.50 7.78 5.65
GPT-5 (fast) 4.11 4.75 5.84 7.21 1.05 3.13 7.30 4.77
LLMs+Reasoning
Qwen3-235B (think) 5.00 5.33 6.64 7.88 1.12 3.90 7.38 5.32
Gemini 2.5 Pro (think) 4.88 5.81 6.99 8.09 2.23 4.40 7.73 5.73
Claude Opus 4.5 (think) 4.96 5.48 6.68 8.10 2.27 4.22 7.73 5.63
GPT-5 (think) 5.57 6.18 7.00 8.06 2.11 4.16 8.08 5.88
LLMs+Reasoning+WebSearch
Qwen3-235B (think+search) 4.05 4.34 5.68 6.83 5.22 5.45 7.06 5.52
Claude Opus 4.5 (think+search) 4.52 5.13 5.99 7.41 7.03 7.62 7.37 6.44
GPT-5 (think+search) 5.57 6.08 6.97 8.15 5.63 6.17 8.11 6.67
Deep Research
WebThinker 4.11 4.64 5.51 7.35 6.21 6.40 7.13 5.91
Qwen3-235B (deep) 4.13 4.69 4.85 7.06 6.55 7.90 7.43 6.09
Gemini 2.5 Pro (deep) 4.71 5.37 6.25 7.59 6.01 7.61 7.39 6.42
Claude Opus 4.5 (deep) 4.53 5.22 5.69 7.22 6.04 5.66 7.57 5.99
OpenAI (deep) 4.67 5.29 6.28 7.66 7.14 6.89 7.48 6.49
Table 1: Baseline models’ expert report evaluation results (updated). In each column, the highest score is marked in bold and the second highest score is underlined.